Explain and show the steps in multiplying and dividing the following:
(2x + 3)(2x - 3)

Answer by wuwei96815(245)

Every term in the second parentheses must be multiplied by every term in the first parentheses. Then, you add all like terms. Example:
2x times (2x-3) = 4x^2 -6x
3 times (2x-3) = 6x-9
4x^2-6x+6x-9 = 4x^2-9